# Benchmarking Resilience and Sensitivity of Polyurethane-Based Vision-Based Tactile Sensors

### v1

@@ III-A3 Abrasion
To test the gel’s abrasion resistance, we abrade it with a custom setup shown in Fig. 2D. We wrap 150 grit sandpaper around a 3D printed 95.5 mm diameter wheel controlled with a brushed DC motor. The robot arm presses the gel against the sandpaper with 5 N of force. We then spin the wheel against the gel at constant velocity for 8 m in 2 m increments, where an unloaded sensor image is captured after each increment. The sandpaper is replaced before every full 8 m trial and used for its entire duration.
@@ III-A4 Evaluation
We evaluate all resilience tests with the same metric. For each cycle, we record 10 RGB frames of both the unloaded and loaded sensor readings at 30 frames per second with QVGA resolution (320x240 pixels) and average them to get one unloaded and one loaded image per cycle. For abrasion, we only record unloaded frames. To evaluate sensor damage, we calculate the mean absolute error ($MAE$) between the image at the current cycle and the image at the first cycle separately for unloaded and loaded images:
@@ III-B Sensitivity
Improving sensor resilience typically reduces sensitivity. To capture these downsides, we propose two tests to evaluate force sensitivity and spatial sensitivity.
@@ III-B1 Force Sensitivity
We evaluate the force sensitivity of our different sensor gels, without durability testing, by loading them up to 40 N on a 4 mm spherical tip indenter at a rate of 2e-6 m/s. We use the same setup used for cyclic compression (Fig. 2A). We restrict the maximum load to 40 N due to hardware limitations, although this range is sufficient for distinguishing silicone and polyurethane performance. During a single loading and unloading, we record 320x240 RGB frames at 30 frames per second with the DIGIT sensor and normal forces with a force torque sensor (ATI Axia80-M8). We then calculate the $MAE$ of each frame with respect to the first frame to quantify how much the captured image changes with loading.
@@ III-B2 Spatial Sensitivity
Our approach for assessing spatial sensitivity in vision-based tactile sensors is a novel, model-free method that analyzes features directly rather than relying on model training. We accomplish this by performing a frequency-domain analysis of the sensor’s measurement of a periodic, ridged surface. By varying the period of the ridges, we can assess x/y sensitivity; similarly, we can evaluate z-sensitivity by varying amplitude. The process is depicted in Fig. 3.
@@
We then preprocess the images to clean the signal. First, we perform background subtraction between the unloaded and loaded frames, $I_{delta}=I_{loaded}-I_{unloaded}$, to isolate the changes observed due to sensor gel deformation. Next, the image is spatially filtered to remove low-frequency artifacts and high-frequency noise using a difference of Gaussians method. Finally, the image is cropped to limit effects caused by camera distortion and the edges of the gel.
To quantify the sensor’s spatial response, we perform a series of one-dimensional Fast Fourier Transforms across each row of pixels spanning the ridged pattern for each color channel, centering to eliminate the DC component. We then calculate power spectrums and average across all rows and channels to obtain a single non-normalized power spectral density (PSD) for the sensor’s signal. We use this spectrum to calculate spatial sensitivity using a Signal-to-Noise Ratio (SNR). The noise floor is obtained by recording the PSD on a flat, ridgeless surface. Next, we calculate signal power $P_{signal}$ for a ridged surface as the summed power of the two frequency bins closest to the the known ridge frequency. We define the noise power $P_{noise}$ as the summed power of the corresponding frequency bins from the spectrum of the flat surface (Fig. 3). Finally, we calculate SNR in decibels (dB) as follows:
@@
In this study, we test across 10 ridge amplitudes with constant period, and 10 ridge periods with constant amplitude. All ridges were printed with custom settings on a Bambu X1E. We test across amplitudes of 0.005 to 0.05 mm, and across periods of 0.6 mm to 1.5 mm. Ridges of varying magnitude all have the maximum period of 1.5 mm, while ridges of varying period have the same maximum amplitude of 0.6 mm. We selected these ranges experimentally to to capture interesting results, ensuring that both materials could generate a meaningful signal at the largest period and amplitude. We also test each ridge with a 2 N and 10 N normal load to observe how it affects spatial sensitivity.
@@ IV RESULTS
Figure 4: Results for cyclic compression (A), cyclic shear on an indenter (B), cyclic shear on a flat surface (C), and abrasion (D) tests across three gel samples for silicone (SI) and polyurethane (PU). The first row represents sensor images captured at each cycle after unloading, while the second row shows results from loaded images. The raw sensor images in the third row depict the final unloaded sensor image (after 1000 cycles) from sample 1 of each material. Different sets of gels are used for each test.
@@ IV-A Mechanical Resilience
Resilience testing results are presented in Fig. 4. Due to variations in the sensor gel that can occur during the manufacturing process, we perform our mechanical resilience tests on three different gels for each material, resulting in six total trials per test. Different gels are used for each test, thus 24 gels were fabricated in total for these particular results. For visualization of relative silicone and polyurethane gel wear, we include representative unloaded sensor images after the test concluded from the first of the three gel samples of each material.
@@ IV-A1 Cyclic Compression Loading
Results from the cyclic compressive loading test can be found in Fig. 4A. Silicone samples 1 and 2 fail catastrophically due to puncture within 250-500 cycles, leading to large increases in $MAE$. The third sample undergoes a much smaller increase in MAE throughout the test. This gel’s image transfer layer undergoes permanent visible deformation from the indenter, but puncture does not occur. All three polyurethane gels see minimal change in the unloaded and loaded images throughout the test, consistently outperforming the silicone gels.
@@ IV-A2 Cyclic Shear Loading
Fig. 4B shows the results of cyclic local shear loading with an indenter for the different gel materials. Silicone samples 1 and 2 encounter catastrophic failure with puncture and tearing occurring before 500 cycles of loading. The third sample undergoes a less noticeable over the course of the test. Puncture still occurs for this gel, but the hole is clean and only induces a small increase in $MAE$ relative to the other two silicone samples. Again, the polyurethane gels consistently outperform the silicone gels, seeing minimal changes in reading throughout the test.
For transverse shear on a flat surface, results in Fig. 4C show limited change in reading across the 1000 cycles. The silicone gels undergo varying degrees of minor delamination from the acrylic window, with the first sample seeing the largest amount. Still, the $MAE$ is limited to $\sim$7 units per pixel when unloaded and $\sim$5 when loaded for this gel, which is much less than that observed for cyclic compression and shear failures on the indenter. The polyurethane gels provide relatively constant images throughout the cycles, with no visible delamination present.
@@ IV-A3 Abrasion
Abrasion testing results are shown in Fig. 4D. The three silicone gels experience varying degrees of tearing in the image transfer layer, often occurring within 2 m of abrasion and increasing with more distance. The polyurethane gels also experience noticeable wear during the test, although a limited increase in $MAE$ indicates that the sensor reading does not experience much change. During testing, the polyurethane image transfer layer was observed to wear away as small particles rather than tear from the clear base layer in bulk like their silicone counterparts, leading to this improvement in performance.
@@ IV-B Sensitivity
Figure 5: Results for force sensitivity (A) and spatial sensitivity (B) tests. In (A), the mean absolute error (MAE) is calculated with respect to the unloaded image to quantify signal change across force. In (B), the first and second rows show SNR under 2 N and 10 N loads, respectively. The left and right columns represent tests with varying amplitude (constant period) and period (constant amplitude).
@@ IV-B1 Force Sensitivity
Results in Fig. 5A show that silicone gels provide much higher force sensitivity for loads less than 10 N. However, once the load increases beyond this point, the signal from the silicone begins to saturate. Conversely, the polyurethane gel provides a lower but more consistent sensitivity across the entire loading range, indicating its usefulness for higher load applications.
@@ IV-B2 Spatial Sensitivity
The results from our spatial sensitivity test, shown in Fig. 5B, illustrate the performance of both materials across the different ridged surfaces and loading conditions. With a 2 N load, the silicone gels generally perform as well as or better than their polyurethane counterparts. For constant period tests, both gels demonstrate a gradual increase in performance with increasing amplitude, with SNR plateauing at 10-15 dB for amplitudes larger than 0.03 mm. For constant amplitude tests, silicone gels provide much clearer signals, especially for periods of 1.0 mm or smaller. When the load is increased to 10 N, the SNR increases for polyurethane but reduces for silicone, resulting in better relative performance for polyurethane across most ridges. The reduction of silicone gel performance at higher load can be attributed to an increase in noise power and may not reflect a decrease in the sensor’s ability to detect the ridges. We discuss this further in Section V. Both sensors show the same general trends across surfaces, seeing plateaus in performance once the ridges reach an amplitude of 0.035 mm or a period of 1.1 mm.
@@ V DISCUSSION
Our results show that polyurethane-based vision-based tactile sensor gels can outlast silicone-based gels across multiple types of wear, including local compression and shear, transverse shear across the whole gel, and abrasion. Additionally, the polyurethane gels may provide improved sensitivity for large forces, where silicone gels may saturate or become noisy due to large deformations. We propose that each sensor has its use case. At low forces, we find that silicone provides improved force and spatial sensitivity, making it an ideal choice for tasks requiring higher precision. However, if sensor reliability or operation at high forces are necessary, our results suggest that polyurethane gels offer an advantage.
We note that our evaluation techniques face certain limitations. The $MAE$ used in our resilience tests shows how much the sensor reading has changed without characterizing how this change affects sensor performance. It is possible that performance can be recovered through sensor recalibration or better models, but we leave this to future work. Our resilience tests–while designed to be representative of real wear–are also highly controlled, with set contact geometries and loading forces. Thus, we do not expect the test lifetimes of our sensors to directly represent their cyclic lifetime in application, and we primarily emphasize the relative performance of each material in these tests.
Although not validated in this work, we hypothesize that our spatial sensitivity results correspond with each gel’s minimum detectable geometries, i.e. that a gel’s signal strength for a ridge of certain amplitude and period reflects its ability to detect features of similar size. The strength of this relationship should be explored in future work. However, sensor performance in our evaluation depends also on loading force. Increasing the loading force can improve signal power, leading to improved SNR as seen with the polyurethane gels. But, larger forces also increase the bulk deformation of the sensor, generating large changes in the sensor image that remain after background subtraction. When the sensor is loaded onto a flat surface to define the noise floor, these changes can increase noise power across frequency bins, leading to reduced SNR for the silicone gels. In application, inference models can be trained on high loads directly and would likely not suffer the same performance drop. We acknowledge this as a limitation of our model-free evaluation, and future work should investigate different pre-processing techniques and noise floor definitions to reduce this effect.
In this work, we only compare between one type of polyurethane gel and one type of silicone gel. Different polyurethanes, silicones, and fabrication techniques may result in changes to relative performance and should be explored in future work.
@@ VI DEMONSTRATION
Figure 6: For demonstration, we perform repeated loosening and tightening of a bottle cap with a sensorized gripper (B). A trained CNN predicts two points that define a simplified contact region shown with dashed lines (C). Mean absolute errors (MAE) of model predictions are calculated for the soft and firm grasps performed during tightening at each cycle (A). The silicone gel tears after 16 cycles (D, top) and is unable to carry on, while the polyurethane gel remains undamaged (D, bottom).
To demonstrate the resilience and performance of both sensors in application, we perform a repeated task of loosening and tightening a bottle cap (Fig. 6B), which induces both compressive and shear loads as well as local surface wear. We use a Robotiq 2F-140 parallel jaw gripper attached to a UR-10 robot arm. For each material, we train a convolutional neural networks (3 convolution layers, 1 hidden dense layer) to predict two points defining a simplified contact region (Fig. 6C). Training data includes background-subtracted images with hand annotations from 243 grasps of varying force and location, augmented by random translations with a custom data generator during training.
We then conduct (up to) fifty cycles of the task, first loosening and then re-tightening a ridged bottle cap to a max torque of 0.8 N-m. During the re-tightening phase, the robot performs a small initial rotation with a soft grasp, resets, and then grasps firmly to tighten the remainder of the way, resulting in two grasps per cycle. The grasp force of the Robotiq 2F-140 is not directly controllable, so the force parameters are set experimentally to ensure a secure grasp. The sensor images from both re-tightening grasps (before any torque is applied) are annotated by hand, and then predictions are made with the trained models. The mean absolute error of the model predicted contact region at each cycle is shown for both grasps (soft and firm) and both materials 6A.
The silicone gel fails via tearing (6) after the 16th cycle of tightening, resulting in increased prediction errors for both grasps in the 17th cycle. After failure, the gel does not provide enough friction to twist the cap, resulting in an early stoppage. The polyurethane gel successfully completes all 50 cycles, maintaining consistent performance without visible failure. This demonstration shows an example of a real-life task where polyurethane VBTS gels can provide improved resilience over silicone gels while still providing useful tactile signals.

### v2

@@ III-A3 Abrasion
To test the gel’s abrasion resistance, we abrade it with a custom setup shown in Fig. 2D. We wrap 150 grit sandpaper around a 3D printed 95.5 mm diameter wheel controlled with a brushed DC motor. The robot arm presses the gel against the sandpaper with 5 N of force. We then spin the wheel against the gel at constant velocity for 8 m in 2 m increments, where an unloaded sensor image is captured after each increment. The sandpaper is replaced before every full 8 m trial and used for its entire duration. While this form of abrasion may be extreme compared to actual applications, it provides a standard way to compare sensor resilience for a worst-case representation of wear.
@@ III-A4 Evaluation
We evaluate all resilience tests with the same metric. After each cycle, we record 10 RGB frames of the unloaded sensor reading at 30 frames per second with QVGA resolution (320x240 pixels) and average them to get one image per cycle. To evaluate sensor damage, we calculate the mean absolute error (MAEMAE) between the average image at the current cycle and the average image at the first cycle:
@@ III-B Sensitivity
Improving sensor resilience typically reduces sensitivity. To assess this in our polyurethane gels, we propose two tests to evaluate force sensitivity and spatial sensitivity.
@@ III-B1 Force Sensitivity
We evaluate the force sensitivity of our different sensor gels by loading them up to 40 N on a 4 mm spherical tip indenter at a rate of 2e-6 m/s. We use the same setup used for cyclic compression (Fig. 2A). We restrict the maximum load to 40 N due to hardware limitations, although this range is sufficient for distinguishing sensor performance. During a single loading and unloading, we record 320x240 RGB frames at 30 frames per second with the DIGIT sensor and normal forces with a force torque sensor (ATI Axia80-M8). We then calculate the MAEMAE of each frame with respect to the first frame to quantify how much the captured image changes with loading.
@@ III-B2 Spatial Sensitivity
Our approach for assessing spatial sensitivity in vision-based tactile sensors is a novel, learning-free method that analyzes features directly rather than relying on model training. We accomplish this by performing a frequency-domain analysis of the sensor’s measurement of a periodic, ridged surface. By varying the period of the ridges, we can assess x/y-sensitivity; similarly, we can evaluate z-sensitivity by varying amplitude. The process is depicted in Fig. 3.
@@
We then preprocess the images to clean the signal. First, we perform background subtraction between the unloaded and loaded frames, $\Delta I=I_{loaded}-I_{unloaded}$, to isolate the changes observed due to sensor gel deformation. Next, the image is spatially filtered to remove low-frequency artifacts and high-frequency noise using a difference of Gaussians method. This involves defining two Gaussian kernels: a narrow kernel to smooth high-frequency noise, and a wide kernel to isolate low-frequency artifacts like uneven lighting:
$G(x,y,\sigma)=\frac{1}{2\pi\sigma^{2}}e^{-\frac{x^{2}+y^{2}}{2\sigma^{2}}}$ (2)
where the standard deviation $\sigma$ determines the smoothing extent. To obtain the final filtered image $D(x,y)$, we convolve the delta image $\Delta I(x,y)$ with the difference of the two Gaussian kernels:
$D(x,y)=\Delta I(x,y)*\left(G(x,y,\sigma_{low})-G(x,y,\sigma_{high})\right).$ (3)
Finally, the image is cropped to limit effects caused by camera distortion and the edges of the gel.
To quantify the sensor’s spatial response, we perform a series of one-dimensional Fast Fourier Transforms across each row of pixels spanning the ridged pattern for each color channel, centering to eliminate the DC component. We then calculate power spectrums and average across all rows and channels to obtain a single non-normalized power spectral density (PSD) for the sensor’s signal. We use this spectrum to calculate spatial sensitivity using a Signal-to-Noise Ratio (SNR). The noise floor is obtained by recording the PSD on a flat, ridgeless surface. Next, we calculate signal power $P_{signal}$ for a ridged surface as the summed power of the two frequency bins closest to the known ridge frequency. We define the noise power $P_{noise}$ as the summed power of the corresponding frequency bins from the spectrum of the flat surface (Fig. 3). Finally, we calculate SNR in decibels (dB) as follows:
@@
In this study, we test across 10 ridge amplitudes with constant period, and 10 ridge periods with constant amplitude. All ridges were printed with custom settings on a Bambu X1E. We test across amplitudes of 0.005 to 0.05 mm, and across periods of 0.6 mm to 1.5 mm. Ridges of varying magnitude all have the maximum period of 1.5 mm, while ridges of varying period have the same maximum amplitude of 0.6 mm. We selected these ranges experimentally to capture interesting results, ensuring that all materials could generate a meaningful signal at the largest period and amplitude. The normal force applied between the sensor and surface is also important for the strength of the signal. Intuitively, harder materials will deform less and thus produce a weaker signal compared to softer materials. So, we also test both 2 N and 10 N normal forces across all ridge configurations.
@@ IV RESULTS
Figure 4: Results for cyclic compression (A), cyclic shear on an indenter (B), cyclic shear on a flat surface (C), and abrasion (D) tests across three gel samples for silicone (SI; shown in the top row in red), a 30A hardness polyurethane (PU30; shown in the second row in purple), and a 50A hardness polyurethane (PU50; shown in the third row in cyan). Each color shade represents a different sample, and grey dots represent results from the other materials. The raw sensor images in the fourth row depict the final sensor image (after 1000 cycles) from the worst-performing sample of each material. Different sets of gels are used for each test.
@@ IV-A Mechanical Resilience
Resilience testing results are presented in Fig. 4. Due to the time required for fabrication and the single-use nature of samples for these tests, we limit testing to three different gels for each material and test condition. [^1] To complete resilience testing, 36 unique samples were required, not counting those used for pilot testing. This is not enough for statistical analysis but still provides some idea of sample variance. For qualitative assessment, we include representative sensor images after the test concluded from the worst-case sample of each material.
@@ IV-A1 Cyclic Compression Loading
Results from the cyclic compressive loading test can be found in Fig. 4A. Two of three silicone samples fail catastrophically due to puncture within 250-500 cycles, leading to large increases in MAEMAE. The third sample undergoes a much smaller increase in MAE throughout the test. This gel’s image transfer layer undergoes permanent visible deformation from the indenter, but puncture does not occur. The three 30A hardness polyurethane (PU30) gels see some minor change across the 1000 cycles but perform much better than the SI gels on average. The 50A hardness polyurethane (PU50) gels undergo the least amount of change during the test, consistently outperforming both other materials.
@@ IV-A2 Cyclic Shear Loading
Fig. 4B shows the results of cyclic local shear loading with an indenter for the different gel materials. Two silicone samples encounter catastrophic failure with puncture and tearing occurring before 500 cycles of loading. The third sample is also punctured, but the hole is clean and only induces a small increase in MAEMAE relative to the other two silicone samples. The PU30 gels all experience some degree of failure via tearing of the reflective image transfer layer, showing little to no performance benefit over SI. Again, the PU50 gels consistently outperform the other two materials, seeing minimal changes in reading throughout the test.
For transverse shear on a flat surface, results in Fig. 4C show limited change in reading across the 1000 cycles. The SI gels undergo varying degrees of minor delamination from the acrylic window, with one sample seeing a particularly large amount. Still, the MAEMAE is limited to $\sim$7 units per pixel, which is much less than that observed for cyclic compression and shear failures with the indenter. PU30 slightly outperforms SI on average but still shows some change over time; however, no visible delamination was observed upon inspection. The PU50 gels provide relatively constant images throughout the cycles, with no visible delamination present.
@@ IV-A3 Abrasion
Abrasion testing results are shown in Fig. 4D. The SI gels experience varying degrees of tearing in the image transfer layer, often occurring within 2 m of abrasion and increasing with more distance. PU30 largely shows better resistance to abrasion than SI, although one sample undergoes tearing in the final 2 m and sees a large increase in MAE. PU50 outperforms SI and PU30, showing little change in MAE for all three samples. However, the gels still exhibited obvious signs of wear. During testing, the polyurethane image transfer layer was observed to wear away as small particles rather than tear from the clear base layer in bulk like their silicone counterparts, leading to this improvement in performance.
@@ IV-B Sensitivity
Figure 5: Results for force sensitivity (A) and spatial sensitivity (B) tests. In (A), the mean absolute error (MAE) is calculated with respect to the unloaded image to quantify signal change across force. In (B), the first and second rows show SNR under 2 N and 10 N loads, respectively. The left and right columns represent tests with varying amplitude (constant period) and period (constant amplitude). The diamond markers connected by a solid line represent the mean SNR for each material.
@@ IV-B1 Force Sensitivity
SI provides the highest force sensitivity at low loads (<10<10 N), as indicated by the slope magnitudes in Fig. 5A. However, the gel reaches its maximum deformation at this point, and the signal becomes totally saturated. PU30 shows force sensitivity similar to that of SI and maintains a clean signal up to 20 N but exhibits a large loading/unloading hysteresis band. Anecdotally, we noted that Clear Flex 30 (for PU30) exhibits much greater viscoelasticity than Solaris (SI) and Clear Flex 50 (PU50), which supports these results. The PU50 gel provides a lower but more consistent sensitivity across the entire loading range, indicating its usefulness for higher load applications.
@@ IV-B2 Spatial Sensitivity
The results from our spatial sensitivity test, shown in Fig. 5B, illustrate the performance of the materials across the different ridged surfaces and loading conditions. With a 2 N load, the SI gels generally perform as well as or better than their PU30 and PU50 counterparts. For constant period tests, all gels demonstrate a gradual increase in performance with increasing amplitude, with SNR plateauing at 10-15 dB for amplitudes larger than 0.03 mm. For constant amplitude tests, SI provides much clearer signals, especially for periods of 1.0 mm or smaller. When the load is increased to 10 N, the SNR increases for PU30 and PU50 but reduces for silicone, resulting in better relative performance for polyurethane across most ridges. The reduction of SI performance at higher load can be attributed to an increase in noise power and may not reflect a decrease in the sensor’s ability to detect the ridges. We discuss this further in Section V. Both materials show the same general trends across surfaces, seeing plateaus in performance once the ridges reach an amplitude of 0.035 mm or a period of 1.1 mm. Interestingly, PU50 provides a signal similar to or greater than PU30 across the majority of tests. Similar to SI, we suspect that the lower stiffness of PU30 leads to a stronger noise floor, which reduces the SNR.
@@ V DISCUSSION
Our results show that the polyurethane gels, with higher hardness and better adhesion to acrylic, can outlast silicone-based gels across multiple types of wear, including local compression and shear, transverse shear across the whole gel, and abrasion. Additionally, the polyurethane gels may provide improved sensitivity for large forces, where silicone gels may saturate or become noisy due to large deformations. We propose that each sensor has its use case. At low forces, we find that silicone provides improved force and spatial sensitivity, making it an ideal choice for tasks requiring higher precision. The 30A hardness polyurethane gel shows improved durability over silicone with small reductions in force and spatial sensitivity, making it a useful alternative in precise applications where silicone may fail–although the issue of hysteresis should be investigated further. If reliability is a much larger concern than sensitivity, our results suggest that the 50A hardness polyurethane gel offers an advantage.
We select the silicone gel formulation due to its representativeness of common VBTS solutions (e.g. DIGIT, GelSight) and ready availability, and the polyurethane formulations based on the softest available water-clear polyurethane rubbers. Higher hardness silicones are available from industrial vendors with long lead times and higher prices and are therefore omitted from this study. However, to fully understand the individual benefits of material and hardness changes, future work should perform these tests across a wider range of gel compositions.
We note that our evaluation techniques face certain limitations. Our resilience tests–while designed to be representative of real wear–are also highly controlled, with specific contact geometries and loading forces that are not comprehensive. Thus, we do not expect the test lifetimes of our sensors to directly represent their cyclic lifetime in application, and we primarily emphasize the relative performance of each material in these tests. The MAEMAE used in our resilience tests measures raw sensor change without characterizing how this change affects sensor performance. It is possible that performance can be recovered through sensor recalibration or better models, but we leave this to future work.
Although not validated in this work, we hypothesize that our spatial sensitivity results correspond with each gel’s minimum detectable geometries. However, sensor performance in our evaluation also depends on loading force. Intuitively, higher force should improve spatial sensitivity as the material continues to deform against the surface. But, larger forces also increase the bulk deformation of the sensor, generating large changes in the sensor image that remain after background subtraction. These changes can increase our noise power measurement across frequency bins, leading to reduced SNR for softer materials at high load. This is a limitation of our learning-free evaluation that should be improved upon in future work, as inference models can be trained on high loads directly to avoid such performance drops in application.
@@ VI DEMONSTRATION
Figure 6: (A) For demonstration, we perform repeated loosening and tightening of a bottle cap with a sensorized gripper. (B) A trained CNN predicts two points that define a simplified contact region shown with dashed lines. (C) Gels exhibit varying degrees of damage after completing $i$ cycles. (D) MAE of model predictions are calculated for the soft and firm grasps performed during tightening at each cycle.
To ground our controlled test results in a real application, we perform a bottle cap loosening/tightening task (Fig. 6B), which induces both compressive and shear loads as well as local surface wear. We use a Robotiq 2F-140 parallel jaw gripper with one DIGIT sensor attached to a UR-10 robot arm. To measure sensor performance, we collect 243 contact images of varying force and location and train a small convolutional neural network to predict two hand-annotated points defining a simplified contact region (Fig. 6C).
We then conduct (up to) fifty cycles of the task, first loosening and then re-tightening a ridged bottle cap to a max torque of 0.8 N-m. For variability, retightening is initiated with a soft grasp and finished with a firm grasp. The sensor images from both re-tightening grasps (before any torque is applied) are annotated by hand, and then predictions are made with the trained models.
As shown by the mean absolute error (MAE) of the model predictions, the silicone gel fails via tearing (Fig. 6C) after the 17th cycle of tightening and exhibits an immediate reduction in performance. After failure, the gel does not provide enough friction to twist the cap, resulting in an early stoppage. For the PU30 gel, a tear occurs in the image transfer layer halfway through the test and continues to grow, only impacting performance for soft grasps after 40 cycles. PU50 successfully completes all 50 cycles, maintaining consistent performance without visible failure. This demonstration shows an example of a real-life task where polyurethane VBTS gels can provide improved resilience over silicone gels while still providing useful tactile signals.